Reports 1-1 of 1 Clear search Modify search
takahiro.yamamoto - 19:56 Monday 02 May 2022 (20636) Print this report
Comment to In-air health check of ITMY: larger low frequency noise of H2 geophone (20197)

[Ushiba, YamaT]


- Similar issue on ITMX F3_GAS will be reported as klog#20639
- Glitches on ACC H2 were removed by replacing the ADC board, the ADC adapter board, and the ADC internal cable.
- Glitch problem was solved.

- On the other hand, ADC timing error is occurred with new ADC.
- It shouldn't affect ITMY local measurement.
- Timing error probably affects measurements with channels on other real-time computer.
- This will be discussed in klog#20642.


According to past investigations reported as the thread of klog#20197,
cause of glitches lies on the downstream of AA chassis. But there was no improvement by replacing AA chassis and the SCSI cable between AA chassis and ADC.

Today, we tried to replace stuffs in the IO chassis.
By replacing ADC (S1402710 => S2113341), signal behavior was drastically changed.
Fig.1 shows the time series signals of ADC noise on Geophone channels (unplugged Dsub cable between LVDT Driver chassis and AA chassis).
before replacing (left panels) and after replacing (right panels).

Vertical axes range is unified as 5-count for all plots. We can see the H2 signal was drastically changed.
Other channels (H1, H3, and H4) were also slightly improved. This seems to come from decreasing the amplitude of 60Hz power line.

Also for spectra, H2 signal became same level of other channels as shown in Fig.2,
though H2 signal had been 100 times louder than other signals below 100Hz before replacement.

After connecting Dsub cable from LVDT Driver chassis, we can also see the improvement of noise level below ~50mHz (See also Fig.3).
Note: due to the Earthquake or other reason, suspension resonances were excited in the plot of "after replacing ADC".

As for the glitch issue, we would say it has been solved.

Images attached to this comment
Search Help